OpenStudy (anonymous):

Andy is four years older than Barbara. Carlos is twice as old as Andy. If Barbara is ___?___ years old, then Andy is ___?___ years old and Carlos is ___?___ years old.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

let andy's age be x and barbara's age be y, carlos age be z x=y+4 z=2x u can substitute any value of y to find x and z

OpenStudy (anonymous):

Let Andy be A Barbara be B Carlos be C A=B+4 C=2A

OpenStudy (anonymous):

use trial and error =) example if we let Carlos be 14yrs old therefore Andy is 7, but Andy is 4 years older than Barbara therefore Barbara is 3yrs old. I hope this will help